The average age of 3 people is 15. If the 3rd person is 5 years older than the 1st person and the 2nd person is 4 years younger than the 1st person, what is the age of the 2nd person?
Answer by TimothyLamb(4379)

---
(x + y + z)/3 = 15
z = x + 5
y = x - 4
---
put the system of linear equations into standard form:
---
x + y + z = 45
x - z = -5
x - y = 4
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 14.666667 = 14 + 2/3
y = 10.666667 = 10 + 2/3
z = 19.666667 = 19 + 2/3
---
system matrix
1 1 1
1 0 -1
1 -1 0
---
inverse of system matrix
0.33333333 0.33333333 0.33333333
0.33333333 0.33333333 -0.66666667
0.33333333 -0.66666667 0.33333333
---
determinant of system matrix = -3
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php